# Bibliography.bib

### v1

@@ 1. Introduction
In [Th1, Th2], Thurston established the h-principle for foliations. One of the ideas that goes into the proof is to decompose the manifold using a triangulation whose simplices are in general position with respect to a given starting distribution; see Definition 8.6. To construct such triangulations, Thurston states what is now known as the ‘jiggling lemma’:
Lemma (Thurston’s jiggling lemma [Th2]).
@@
This result has found numerous applications in the construction of geometric structures on manifolds, beyond foliations. These include Haefliger structures with transverse geometry [LM], contact structures [colin1999stabilite, Vo16], and Engel structures [CPPP, PV].
@@ 1.1. Jiggling sections
Let $E\to M$ be a fiber bundle and let $s,s^{\prime}:M\to E$ be piecewise smooth sections with respect to triangulations $T$ and $T^{\prime}$ of $M$, respectively. We say that $(s^{\prime},T^{\prime})$ is an $\varepsilon$-jiggling of $(s,T)$ if $T^{\prime}$ is a subdivision of $T$ and $d_{C^{1}}(s,s^{\prime})<\varepsilon$, where $\varepsilon$ is a positive constant if $M$ is compact, and a positive function otherwise.
@@
We now state our main theorem: {restatable*}thmjigglingMfd Let $M$ be a manifold, possibly open, endowed with a smooth triangulation $T:K\to M$. Consider a fiber bundle $E\to M$ and an open and fiberwise dense differential relation ${\mathcal{R}}\subset J^{1}(E)$. Then, given
@@
  • $s^{\prime}$ is a piecewise solution of ${\mathcal{R}}$ with respect to $T^{\prime}$, and
@@
The idea of jiggling ‘one color at a time’ was inspired by the coloring argument used by Donaldson in [Do], which he used to achieve quantitative transversality in the asymptotically holomorphic setting. Coloring allows us to avoid various size and distance estimates in the Grassmannian appearing in Thurston’s original proof. It is in this simplification that our argument deviates strongly from his, and therefore it can be more readily generalized.
@@ 1.2. Jiggling as an h-principle
This leads us to define the simplicial set ${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})$ of piecewise smooth solutions of ${\mathcal{R}}$ by letting $({\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}))_{n}$ be the set of piecewise smooth sections $s:M\times\Delta^{n}\to E$ that solve ${\mathcal{R}}$. We compare this simplicial set to the singular complex $\operatorname{Sing}(\Gamma^{0}(E))$ of the space $\Gamma^{0}(E)$ of continuous sections of $E$. {restatable*}thmwheSSets Let ${\mathcal{R}}\subset J^{1}(E)$ be an open and fiberwise dense differential relation. The inclusion ${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})\hookrightarrow\operatorname{Sing}(\Gamma^{0}(E))$ is weak homotopy equivalence of simplicial sets. We pose a couple of interesting open questions: Can a version of Theorem 1.2 be established for higher order jets? Moreover, can the assumption of fiberwise density be replaced by some form of ampleness (as is required in convex integration)?
We remark that Theorem 1.2 is an h-principle without homotopical assumptions. Concretely: since our solutions $s:M\rightarrow E$ are meant to be piecewise, their first jet extensions $j^{1}s:M\rightarrow{\mathcal{R}}$ can have discontinuities along the triangulation. As such, they do not define a homotopy class of a map $M\rightarrow{\mathcal{R}}$. This is why no formal data appears in the statement.
The best known example of an h-principle without homotopical assumptions is the arborealization program in symplectic/contact topology [alvarez2020positive, alvarez2021arborealizationI, alvarez2021arborealizationII]. One of its goals, which is similar in flavor to Thurston’s jiggling, is to show that front projections of legendrians can be simplified so that their singularities are of so-called arboreal type. This can then be used to simplify the singularities of skeleta of Weinstein manifolds.
@@ 1.3.1. General position
The condition that a triangulation is in general position with respect to a distribution $\xi$ is, of course, triangulation dependent. Because of this we have to do a bit of work to phrase it using an open and fiberwise dense relation. We do this in Section 8.3.2, establishing Thurston’s jiggling lemma (Corollary 8.10). Something that is new in this direction is that we establish the result for manifolds that are not necessarily compact, which was not addressed by Thurston [Th2]. In fact, from a technical standpoint, an interesting feature of the proof is that we have to develop some theory regarding crystalline subdivision in the non-compact setting, which we do in Section 3.5.
Another application is the construction of triangulations that are in general position with respect to multiple distributions at the same time; a statement that has appeared already in applications, e.g. [Vo16, PV].
@@ 1.3.2. Piecewise maps of full rank
${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}_{\operatorname{transv}})$ $\hookrightarrow\operatorname{Sing}(C^{0}(M,N)),$
${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}_{\operatorname{imm}})$ $\hookrightarrow\operatorname{Sing}(C^{0}(M,N))\text{ if }\dim M<\dim N,\text{ and }$
${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}_{\operatorname{subm}})$ $\hookrightarrow\operatorname{Sing}(C^{0}(M,N))\text{ if }\dim M>\dim N.$
@@ 1.3.3. Piecewise differentiable geometric structures
Definition 1.2 tells us that every non-vanishing 1-form is homotopic to a piecewise smooth contact form. Denoting by ${\mathcal{R}}_{\operatorname{cont}}$ the relation for contact forms, we obtain the following h-principle: {restatable*}corappContactWhe The inclusion ${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}_{\operatorname{cont}})\hookrightarrow\operatorname{Sing}\left(\Gamma^{0}\left(\Lambda^{1}\left(T^{*}M\right)\setminus\{0\}\right)\right)$ is a weak homotopy equivalence.
This complete flexibility in the piecewise smooth setting should be compared to the rigidity that appears in the study of other geometric structures of low regularity. Examples include $C^{0}$-contact structures [MuSp, El97wave, RZ20], $C^{0}$-symplectic structures [Op09, HLS15], Lipschitz symplectic structures [Jo22], and PL symplectic structures [BD24].
@@ 1.5. Acknowledgements
The authors want to thank Miguel Barata for reading a draft of this paper.
@@ 2.1. Jet spaces and h-principles
In this section we briefly recall the standard terminology on (first order) jet spaces and h-principles. For a more elaborate overview we refer to [CiElMi, Gr86].
@@
which allows us to compare the space of solutions ${\operatorname{Sol}}({\mathcal{R}})$ to the space of formal solutions ${\operatorname{FSol}}({\mathcal{R}})$, which is more algebraic in nature. The theory of h-principles concerns itself with the question whether the above inclusion is a weak homotopy equivalence. If it is, we say that ${\mathcal{R}}$ satisfies the (complete) h-principle. Gromov’s method of flexible sheaves [gromov1969stable] shows that the h-principle holds for a large class of relations (open and ${\operatorname{Diff}}$-invariant) on open manifolds. Partial and negative results have also been proven, such as [Be, El92, taubes1994seiberg].
@@
  • relatively if given a section $s\in{\operatorname{FSol}}({\mathcal{R}})$ such that $s|_{C}$ is holonomic, there exists a holonomic section $s^{\prime}\in{\operatorname{Sol}}({\mathcal{R}})$, which is homotopic to $s$ relative ${\mathcal{O}p}(C)$, for every closed set $C\subset M$.
  • parametrically if given a continuous family $S:P\to{\operatorname{FSol}}({\mathcal{R}})$, there exists a continuous family $S^{\prime}:P\to{\operatorname{Sol}}({\mathcal{R}})$ homotopic to $S$. Here $P$ is a compact manifold acting as parameter space.
@@ 2.2.1. Simplicial complexes
  • if $\sigma\in K$ then its face is also in $K$, and
  • if $\sigma,\sigma^{\prime}\in K$ then $\sigma\cap\sigma^{\prime}$ is either empty or an element of $K$.
We denote by $K^{(\mathrm{top})}$ the set of top dimensional simplices of a simplicial complex $K$, and can also define subcomplexes in an obvious manner.
@@ 2.2.3. Subdivisions and isomorphisms
A more “local” version for two simplicial complexes is the following:
@@
  • $K_{1}$ and $K_{2}$ are isomorphic, and
  • $Q=|K_{1}|=|K_{2}|$.
@@ 2.2.4. Piecewise linear maps
If we want to emphasize the role of $K$, we will write that $(f,K)$ is a piecewise linear map or that $f$ is piecewise linear with respect to $K$. We denote the set of piecewise linear maps by $C_{\operatorname{PL}}(P,{\mathbb{R}}^{n})$. We topologize it below.
@@ 2.2.5. Piecewise smooth maps
The set of piecewise smooth maps will be denoted by $C_{\operatorname{PS}}(P,N)$. We note that $C_{\operatorname{PS}}(P,{\mathbb{R}}^{n})$ contains $C_{\operatorname{PL}}(P,{\mathbb{R}}^{n})$. The set $C_{\operatorname{PS}}(P,N)$ can readily be endowed with the $C^{0}$-topology (either the weak or the strong version), by interpreting it as a subset of $C^{0}(P,N)$.
To define the $C^{r}$-topology in $C_{\operatorname{PS}}(P,N)$, it is convenient to assume that $P$ is of pure dimension. We fix an auxiliary triangulation $K$ of $P$ to make sense of $J^{r}(\Delta,N)$ for each top simplex $\Delta\in P$. On each $J^{r}(\Delta,N)$ we fix a metric, so that we can make sense of the $C^{r}$-metric on $C_{\operatorname{PS}}(\Delta,N)$. Then we define:
@@
Consider a pair of maps $f_{1},f_{2}\in C_{\operatorname{PS}}(P,N)$, each piecewise smooth with respect to a triangulation $K_{i}$ of $P$, for $i=1,2$. Consider a triangulation $K^{\prime}$ subdividing $K_{1}$, $K_{2}$ and $K$. We write $d_{C^{r}}(f_{1},f_{2})<\varepsilon$ if for every top dimensional simplex $\Delta\in K^{\prime}$ we have $d_{C^{r}}(f_{1}|_{\Delta},f_{2}|_{\Delta})<\varepsilon$. Here we let $\varepsilon$ be a positive constant if $P$ is compact, and a positive function $P\to{\mathbb{R}}_{+}$ if $P$ is not compact.
We observe that the above distance on $C_{\operatorname{PS}}(P,N)$ depends on our choice of metrics on each $J^{r}(\Delta,N)$ and therefore it also depends on $K$. It does however not depend on the choice of $K^{\prime}$. The underlying topologies (known as respectively the weak and strong) do not depend on any of these choices.
@@
Consider a pair of maps $f_{1},f_{2}\in C_{\operatorname{PS}}(P,N)$, each compatible with a triangulation $K_{i}$ of $P$, for $i=1,2$. Assume that the simplicial complex $K_{2}$ subdivides $K_{1}$, and that $d_{C^{1}}(f_{1},f_{2})<\varepsilon$. Then the map $(f_{2},K_{2})$ is an $\varepsilon$-jiggling of $(f_{1},K_{1})$.
@@ 2.2.7. Ordered simplicial complexes
Observe that, in an ordered simplicial complex, each simplex can be uniquely parametrized by the standard simplex in a manner that is affine and order preserving (where the vertices of the standard simplex are ordered using the standard basis in ${\mathbb{R}}^{m}$).
@@ 2.2.8. Colored simplicial complexes
Let $K$ be a simplicial complex of pure dimension. A $C$-coloring of $K$ is a map

### v2

# Jiggling: an h-principle without homotopical assumptions

@@ 1. Introduction
In [30, 31], Thurston established the h-principle for foliations. One of the ideas that goes into the proof is to decompose the manifold using a triangulation whose simplices are in general position with respect to a given starting distribution; see Definition 8.6. To construct such triangulations, Thurston states what is now known as the ‘jiggling lemma’:
Lemma (Thurston’s jiggling lemma [31]).
@@
This result has found numerous applications in the construction of geometric structures on manifolds, beyond foliations. These include Haefliger structures with transverse geometry [20], contact structures [8, 32], and Engel structures [6, 26].
@@ 1.1. Jiggling sections
Consider a fiber bundle $E\to M$ and let $s,s^{\prime}:M\to E$ be piecewise smooth sections with respect to triangulations $T$ and $T^{\prime}$ of $M$, respectively. We say that $(s^{\prime},T^{\prime})$ is an $\varepsilon$-jiggling of $(s,T)$ if $T^{\prime}$ is a subdivision of $T$ and $d_{C^{1}}(s,s^{\prime})<\varepsilon$, where $\varepsilon$ is a positive constant if $M$ is compact, and a positive function otherwise.
@@
We now state our main theorem: {restatable*}thmjigglingMfd Let $M$ be a manifold, possibly open, endowed with a smooth triangulation $T:|K|\to M$. Consider a fiber bundle $E\to M$ and an open and fiberwise dense differential relation ${\mathcal{R}}\subset J^{1}(E)$. Then, given
@@
  • $s^{\prime}$ is a piecewise smooth solution of $\mathcal{R}$} with respect to $T^{\prime}$, and
@@
The idea of jiggling ‘one color at a time’ was inspired by the coloring argument used by Donaldson in [9], which he used to achieve quantitative transversality in the asymptotically holomorphic setting. Coloring allows us to avoid various size and distance estimates in the Grassmannian appearing in Thurston’s original proof. It is in this simplification that our argument deviates strongly from his, and therefore it can be more readily generalized.
@@ 1.2. Jiggling as an h-principle
This leads us to define the simplicial set ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$ of piecewise smooth solutions of $\mathcal{R}$} by letting $({\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}))_{n}$ be the set of piecewise smooth sections $s:M\times\Delta^{n}\to E$ that solve $\mathcal{R}$}. We compare this simplicial set to the singular complex $\operatorname{Sing}(\Gamma^{0}(E))$ of the space $\Gamma^{0}(E)$ of continuous sections of $E$. {restatable*}thmwheSSets Let ${\mathcal{R}}\subset J^{1}(E)$ be an open and fiberwise dense differential relation. The inclusion ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})\hookrightarrow\operatorname{Sing}(\Gamma^{0}(E))$ is weak homotopy equivalence of simplicial sets. We pose a couple of interesting open questions: Can a version of Theorem 1.2 be established for higher order jets? Moreover, can the assumption of fiberwise density be replaced by some form of ampleness (as is required in convex integration)?
We remark that Theorem 1.2 is an h-principle without homotopical assumptions. Concretely: since our solutions $s:M\rightarrow E$ are meant to be piecewise smooth, their first jet extensions $j^{1}s:M\rightarrow{\mathcal{R}}$ can have discontinuities along the triangulation. As such, they do not define a homotopy class of a map $M\rightarrow{\mathcal{R}}$. This is why no formal data appears in the statement.
The best known example of an h-principle without homotopical assumptions is the arborealization program in symplectic/contact topology [3, 1, 2]. One of its goals, which is similar in flavor to Thurston’s jiggling, is to show that front projections of legendrians can be simplified so that their singularities are of so-called arboreal type. This can then be used to simplify the singularities of skeleta of Weinstein manifolds.
@@ 1.3.1. General position
The condition that a triangulation is in general position with respect to a distribution $\xi$ is, of course, triangulation dependent. Because of this we have to do a bit of work to phrase it using an open and fiberwise dense relation. We do this in Section 8.3.2, establishing Thurston’s jiggling lemma (Corollary 8.10). Something that is new in this direction is that we establish the result for manifolds that are not necessarily compact, which was not addressed by Thurston [31]. In fact, from a technical standpoint, an interesting feature of the proof is that we have to develop some theory regarding crystalline subdivision in the non-compact setting, which we do in Section 3.5.
Another application is the construction of triangulations that are in general position with respect to multiple distributions at the same time; a statement that has appeared already in applications, e.g. [32, 26].
@@ 1.3.2. Piecewise maps of full rank
${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}_{\operatorname{transv}})$ $\hookrightarrow\operatorname{Sing}(C^{0}(M,N)),$
${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}_{\operatorname{imm}})$ $\hookrightarrow\operatorname{Sing}(C^{0}(M,N))\text{ if }\dim M<\dim N,\text{ and }$
${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}_{\operatorname{subm}})$ $\hookrightarrow\operatorname{Sing}(C^{0}(M,N))\text{ if }\dim M>\dim N.$
@@ 1.3.3. Piecewise differentiable geometric structures
Definition 1.2 tells us that every non-vanishing 1-form is homotopic to a piecewise smooth contact form. Denoting by ${\mathcal{R}}_{\operatorname{cont}}$ the relation for contact forms, we obtain the following h-principle: {restatable*}corappContactWhe The inclusion ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}_{\operatorname{cont}})\hookrightarrow\operatorname{Sing}\left(\Gamma^{0}\left(\Lambda^{1}\left(T^{*}M\right)\setminus\{0\}\right)\right)$ is a weak homotopy equivalence.
This complete flexibility in the piecewise smooth setting should be compared to the rigidity that appears in the study of other geometric structures of low regularity. Examples include $C^{0}$-contact structures [24, 11, 27], $C^{0}$-symplectic structures [25, 17], Lipschitz symplectic structures [19], and PL symplectic structures [5].
@@ 1.5. Acknowledgements
The authors want to thank Miguel Barata for reading a draft of this paper. This paper forms a part of the first author’s PhD thesis. Hence, the authors would like to thank the reading committee, consisting of Mélanie Bertelson, Kai Cieliebak, Gaël Meigniez, Ieke Moerdijk and Thomas Rot, for reviewing the thesis and offering their feedback.
@@ 2.1. Jet spaces and h-principles
In this section we briefly recall the standard terminology on (first order) jet spaces and h-principles. For a more elaborate overview we refer to [7, 14].
@@
which allows us to compare the space of solutions ${\operatorname{Sol}}({\mathcal{R}})$ to the space of formal solutions ${\operatorname{FSol}}({\mathcal{R}})$, which is more algebraic in nature. The theory of h-principles concerns itself with the question whether the above inclusion is a weak homotopy equivalence. If it is, we say that $\mathcal{R}$} satisfies the (complete) h-principle. Gromov’s method of flexible sheaves [15] shows that the h-principle holds for a large class of relations (open and $\operatorname{Diff}$}-invariant) on open manifolds. Partial and negative results have also been proven, such as [4, 13, 29].
@@
  • relatively if given a section $s\in{\operatorname{FSol}}({\mathcal{R}})$ such that $s|_{{\mathcal{O}p}(C)}$ is holonomic, there exists a holonomic section $s^{\prime}\in{\operatorname{Sol}}({\mathcal{R}})$, which is homotopic to $s$ relative $C$, for every closed set $C\subset M$.
  • parametrically if given a continuous family $S:K\to{\operatorname{FSol}}({\mathcal{R}})$, there exists a continuous family $S^{\prime}:K\to{\operatorname{Sol}}({\mathcal{R}})$ homotopic to $S$. Here $K$ is a compact manifold acting as parameter space.
@@ 2.2.1. Simplicial complexes
  • if $\sigma\in K$ then its faces are also in $K$, and
  • if $\sigma,\sigma^{\prime}\in K$ then $\sigma\cap\sigma^{\prime}$ is either empty or a face of both $\sigma$ and $\sigma^{\prime}$.
We denote by $K^{(\mathrm{top})}$ the set of top-dimensional simplices of a simplicial complex $K$, and can also define subcomplexes in an obvious manner.
@@ 2.2.3. Subdivisions and isomorphisms
A more “local” version of equivalence for two simplicial complexes is the following:
@@
  • $K_{1}$ and $K_{2}$ are isomorphic,
  • $K_{1}$ is the largest subcomplex of $K$ such that $|K_{1}|\subset Q$, and
  • $K_{2}$ is the largest subcomplex of $K^{\prime}$ such that $|K_{2}|\subset Q$.
@@ 2.2.4. Piecewise linear maps
If we want to emphasize the role of $K$, we will write that $(f,K)$ is a piecewise linear map or that $f$ is piecewise linear with respect to $K$. We denote the set of piecewise linear maps by $C_{\mathrm{PL}}(P,{\mathbb{R}}^{n})$. We topologize it below.
@@ 2.2.5. Piecewise smooth maps
The set of piecewise smooth maps will be denoted by $C_{\mathrm{PS}}(P,N)$. We note that $C_{\mathrm{PS}}(P,{\mathbb{R}}^{n})$ contains $C_{\mathrm{PL}}(P,{\mathbb{R}}^{n})$. The set $C_{\mathrm{PS}}(P,N)$ can readily be endowed with the $C^{0}$-topology (either the weak or the strong version), by interpreting it as a subset of $C^{0}(P,N)$.
To define the $C^{r}$-topology in $C_{\mathrm{PS}}(P,N)$, it is convenient to assume that $P$ is of pure dimension. We fix an auxiliary triangulation $K$ of $P$ to make sense of $J^{r}(\Delta,N)$ for each top simplex $\Delta\in P$. On each $J^{r}(\Delta,N)$ we fix a metric, so that we can make sense of the $C^{r}$-metric on $C_{\mathrm{PS}}(\Delta,N)$. Then we define:
@@
Consider a pair of maps $f_{1},f_{2}\in C_{\mathrm{PS}}(P,N)$, each piecewise smooth with respect to a triangulation $K_{i}$ of $P$, for $i=1,2$. Consider a triangulation $K^{\prime}$ subdividing $K_{1}$, $K_{2}$ and $K$. We write $d_{C^{r}}(f_{1},f_{2})<\varepsilon$ if for every top-dimensional simplex $\Delta\in K^{\prime}$ we have $d_{C^{r}}(f_{1}|_{\Delta},f_{2}|_{\Delta})<\varepsilon$. Here we let $\varepsilon$ be a positive constant if $P$ is compact, and a positive function $P\to{\mathbb{R}}_{+}$ if $P$ is not compact.
We observe that the above distance on $C_{\mathrm{PS}}(P,N)$ depends on our choice of metrics on each $J^{r}(\Delta,N)$ and therefore it also depends on $K$. It does however not depend on the choice of $K^{\prime}$. The underlying topologies (known as respectively the weak and strong) do not depend on any of these choices.
@@
Consider a pair of maps $f_{1},f_{2}\in C_{\mathrm{PS}}(P,N)$, each compatible with a triangulation $K_{i}$ of $P$, for $i=1,2$. Assume that the simplicial complex $K_{2}$ subdivides $K_{1}$, and that $d_{C^{1}}(f_{1},f_{2})<\varepsilon$. Then the map $(f_{2},K_{2})$ is an $\varepsilon$-jiggling of $(f_{1},K_{1})$.
@@ 2.2.7. Ordered simplicial complexes
We observe that, in an ordered simplicial complex, each simplex can be uniquely parametrized by the standard simplex in a manner that is affine and order preserving (where the vertices of the standard simplex are ordered using the standard basis in ${\mathbb{R}}^{m}$). Hence to each ordered linear simplex we associate two linear maps. To formalize this, we first introduce some notation.
Notation 2.15.
We denote the affine span of a subset $S\subset{\mathbb{R}}^{n}$ by $\operatorname{ASpan}(S)$. When considering a $d$-simplex $\Delta$, the plane $\operatorname{ASpan}(\Delta)$ is the unique affine $d$-plane containing $\Delta$. Similarly, we denote by $\operatorname{LSpan}(S)$ the linear span of a subset $S\subset{\mathbb{R}}^{n}$.
Definition 2.16.
A linear, ordered $m$-simplex $\Delta=\langle v_{0},\dots,v_{m}\rangle$ in ${\mathbb{R}}^{N}$ defines the unique linear map $T_{\Delta}:{\mathbb{R}}^{m}\to{\mathbb{R}}^{N}$ sending the $i$th basis element $e_{i}$ to $v_{i}-v_{0}$. Translating $\Delta$ by $v_{0}$ provides us with the set $\{-v_{0}\}+\Delta$, and by considering its linear span we obtain the linear map $T_{\Delta}^{-1}:\operatorname{LSpan}(\{-v_{0}\}+\Delta)\to{\mathbb{R}}^{m}$.
@@ 2.2.8. Colored simplicial complexes
Let $K$ be a simplicial complex of pure dimension. A $C$-coloring of $K$ is a surjective map